A loan of $25, 000 is paid off in semi-annual payments over a four year period.   Interest is 6.2% compounded quarterly. What is the size of the payment made at the end of every six months ?

(Please use the financial calculator method (BAII) by showing inputs in the financial calculator)
Answer #1

Semiannual rate = (1 + 0.062 / 4)^2 - 1 = 1.03124 - 1 = 3.124%

PV = -25000

FV = 0

N = 8

I/Y = 0.03124

CPT PMT = 3,580.07

The size of the payment made at the end of every six months = $3,580.07